OpenStudy (anonymous):

x^2/2 + 4/x when is it concave up?

OpenStudy (loser66):

you first! hehehe

OpenStudy (anonymous):

ha so i tried finding second derivative but got different answers than the answer choices so i figured i would start over

OpenStudy (loser66):

ok, show me steps

OpenStudy (anonymous):

should i do quotient rule?

OpenStudy (anonymous):

sooo do i use quotient rule on each term

OpenStudy (anonymous):

ohhhh you can pull out the coefficient

OpenStudy (anonymous):

so f'=

OpenStudy (anonymous):



OpenStudy (loser66):

\[(\dfrac{x^2}{2}+\dfrac{4}{x})' = x -\dfrac{4}{x^2}\]

OpenStudy (anonymous):

correct. i typed it wrong

OpenStudy (loser66):

\[(x -\dfrac{4}{x^2})' = 1+ \dfrac{8}{x^3}\]

OpenStudy (loser66):

now, make it the same denominator, you have f"(x) = \(\dfrac{x^3+8}{x^3}\)

OpenStudy (anonymous):

ok \

OpenStudy (loser66):

for the numerator, you have x = -2 , and 2 complex roots, but ignore the complex roots, so, so far you have x =-2 make f"(x) = 0 ok?

OpenStudy (anonymous):

ya

OpenStudy (loser66):

and with x >-2 the numerator >0 ok?

OpenStudy (loser66):

just numerator, not the whole thing yet, Please, be patient

OpenStudy (anonymous):

so then the denom has a critical number of 0

OpenStudy (loser66):

stop

OpenStudy (anonymous):

yes sir

OpenStudy (loser66):

you should understand what I am saying first

OpenStudy (anonymous):

i do understand that

OpenStudy (loser66):

with x <-2 the numerator <0 OK?

OpenStudy (loser66):